OpenStudy (anonymous):

If h(x) = 2x + 1, find the value of x if h(x) = 11. 5 23 6 10

jhonyy9 (jhonyy9):

2x+1=11 so can you continue ?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

IS IT 10

jhonyy9 (jhonyy9):

subtract from both sides 1 than will be 2x+1-1=11-1 2x=10 divide both sides by 2 x=10/2 =5 x=5 ok ?

OpenStudy (anonymous):

HOW DID YU GET 2X+1=11

jhonyy9 (jhonyy9):

so first there is h(x)=2x+1 and secondly h(x)=11 yes ?

jhonyy9 (jhonyy9):

so because h(x) need to be equal h(x) so than resulted that 2x+1=11

jhonyy9 (jhonyy9):

ok ?